OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 6x - 5y = 4 9x + 4y = -17

OpenStudy (shubhamsrg):

homework ? o.O

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

6x - 5y = 4 (1) 9x + 4y = -17 (2) multiply equation (1) by 3 and equation (2) by 2 and add the 2 equations to eliminate x 18x - 15y = 12 -18x - 8y = 34 --------------- -23y = 46 divide by -23 gives y = -2 substitute into equation 1 to find x 6x - 4(-2) = 4 just evaluate

OpenStudy (campbell_st):

oops equation 2 by -2 and add

OpenStudy (anonymous):

sorry i really suck at math what do you mean evaluate